Determine the length of curve π‘₯ =

2/3

(𝑦 βˆ’ 1)raise to 3/2 when 𝑦 ∈ [1,4]


1
Expert's answer
2021-04-20T07:33:44-0400

"\displaystyle
x = \frac{2}{3} \left(y - 1\right)^{\frac{3}{2}}\\

\frac{\mathrm{d}x}{\mathrm{d}y} = \left(y - 1\right)^{\frac{1}{2}}\\


\left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^2 = y - 1\\


1 + \left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^2 = y\\

\sqrt{1 + \left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^2} = \sqrt{y}\\


\begin{aligned}
\int_1^4\sqrt{1 + \left(\frac{\mathrm{d}x}{\mathrm{d}y}\right)^2}\,\mathrm{d}y &= \int_1^4 \,\,\sqrt{y}\,\mathrm{d}y 
\\&= \frac{2y^{\frac{3}{2}}}{3}\biggr\vert_1^4 = \frac{16 - 2}{3} = \frac{14}{3}
\end{aligned}"